# From LLMs to LLM-based Agents for Software Engineering: A Survey of Current, Challenges and Future

### v1

@@ IX-C Analysis
Figure 8: Illustration of Comparison Framework Between LLM-based Agent [46] and LLM [152] in Software Security and Maintenance
@@ IX-E Evaluation Metrics
The evaluation metrics for the llm in the software security and maintenance are quite diverse. Researchers need to consider various factors such as coverage, efficiency, and reliability of the model or framework. Evaluation Metrics like success rate and pass rate are directly related to the performance of LLMs in different scenarios. In Table X, common standards such as success rate and change rate are frequently used to evaluate the robustness of models when faced with diverse inputs. Time overhead and query number are used to assess the efficiency and resource consumption of models when performing specific tasks. Additionally, ROC AUC, F1 score, and accuracy are important for evaluating the model’s ability to identify vulnerabilities, especially in binary classification tasks. In code repair tasks, metrics such as compilability and plausibility are very common, these metrics ensure that the generated solutions are correct and deployable. Common standards like BLEU and CodeBLEU are used to evaluate the quality and human-likeness of generated code, which helps determine if the model’s capabilities and performance are comparable to human performance. Furthermore, domain-specific metrics like tree edit distance and test pass rate are used to evaluate the effectiveness of LLM applications in specialized fields of software engineering, these metrics are used to address the limitations posed by software security and maintenance. In contrast, while LLM-based agents use evaluation metrics similar to those used for LLMs, such as success rate, they also incorporate more subjective metrics for evaluation. These include appropriateness, relevance, and adequacy, which are human-judged standards. Overall, the evaluation metrics used by agents tend to be simpler and easier to understand than those used for LLMs. This is likely because agents handle high-level tasks, such as the success rate of generating potential vulnerabilities and the frequency of agents calling external tools, so they also need to consider computational and time overheads of the overall architecture.
@@
TABLE X: Evaluation Metrics in Software Security and Maintenance
@@
Vulnerable Virtual Machine
@@ 13 Smart Contract Bugs
1734 Positive Samples,
1810 Negative Samples
F1 score, Accuracy, Consistency
@@
13 Android Applications
from GitHub
Recall, Precision, Correct, Over-fitting, Correct@k
Yes
[164]
@@ 13 Android Applications
15 Vulnerabilities from CVE Lib Success Rate Yes
@@ X-A Experiment Models
In section 3-8, we reviewed and introduced the research on LLMs and LLM-based agent applications in software engineering in recent years. These studies have different research directions and we divided them into six subtopics for classification and discussion. With the advancement of large language models, thousands of models have appeared in the public eye, in order to more intuitively understand the application of large language models in various fields and the use of large language models as the core of intelligent agents, we summarized a total of 117 papers, mainly to discuss the frequency of use of LLMs in the field of software engineering.
@@
Based on the review of 117 papers, our primary focus is on the models or frameworks utilized by the authors in their experiments. This is due to the fact that these papers often include tests of model performance in specific domains, such as evaluating LLaMA’s performance in code generation. Therefore, during our data collection process, we also included models used for comparison purposes, as these models often represent the state-of-the-art capabilities in their respective fields at the time of the study. In summary, across the 117 papers, we identified a total of 79 unique large language models. We visualized the frequency of these model names in a word cloud for a more intuitive representation, as shown in Figure.9. From the figure, we can observe that models such as GPT-3.5, GPT-4, LLaMA2, and Codex are frequently used. Although close source LLMs cannot be locally deployed or further trained, their exceptional capabilities make them a popular choice for comparison in experiments or for data augmentation, where GPT-4 is used to generate additional data to support the research model frameworks. For instance, researchers might use OpenAI’s API to generate initial text and then employ locally deployed models for further processing and optimization [76] [122] [119] [113].
@@
Therefore, it is not difficult to see that the use of general large models with superior performance to assist development or as a measurement standard has been increasingly used in the vertical field of software engineering in the past two years. In addition, for some fields that have never been touched by LLMs before, many researchers first refer to the model ChatGPT and conduct various performance experiments on the newer GPT-4 [55] [58] [64]. Those models can be integrated into larger systems and combining with other machine learning models and tools, these models can be used to generate natural language responses, while another model handles intent recognition and dialogue management.
Figure 9: Experiment Models Usage WordCloud
@@
Although the word cloud provides a rough overview of model usage frequency, it lacks detailed information. To gain deeper insights, we combined a grouped bar chart and a stacked bar chart to further analyze the usage of models in studies across different subtopics. The corresponding bar charts are presented in Figure.10. During the analysis, we found that a large number of models appeared only once. Including these in the bar chart would have made the overall representation cluttered. Therefore, we excluded models that appeared only once and focused on the versatility of the remaining models. On the left side of each subtopic, we depict the models used in LLM-related studies, with the models used in LLM-based agent-related studies highlighted in red-bordered bars. From the figure, it is evident that in the Autonomous Learning and Decision Making subtopic, the number of models used in LLM-based agent-related studies is quite high. Specifically, GPT-4 and GPT-3.5 were used in 10 out of 18 papers and 15 out of 18 papers, respectively. In this subtopic, studies commonly utilized GPT-3.5/4 and LLaMA-2 for research and evaluation. During our analysis, we found that many studies on LLM-based agents evaluated the agents’ ability to mimic human behavior and decision-making or perform some reasoning tasks [103] [111] [108]. Since these studies do not require local deployment, they mainly assess the performance of state-of-the-art models in specific directions, leading to the frequent use of the GPT-family models. Frameworks like [98] [36] constructed LLM-based agents by calling the GPT-4 API, using verbal reinforcement to help language agents learn from their mistakes. Due to the limitations of GPT models, many studies also used LLaMA as the LLM for agents, fine-tuning it on the generated datasets to evaluate the emergence of knowledge and capabilities. Overall, we found that in the Autonomous Learning and Decision Making subtopic, LLM-based agents often use multiple models for testing and performance evaluation in a single task, this results in a significantly higher model usage frequency in this topic compared to others.
@@
Figure 10: Experiment Models Usage in Different Subtopics (REQ DENOTES ”Requirement Engineering and Documentation”, CODE DENOTES ”Code Generation and Software Development”, AUTO DENOTES ”Autonomous Learning and Decision Making”, DES DENOTES ”Software Design and Decision Making”, SEC DENOTES ”Software Security and Maintenance”)
@@
Therefore, researchers and developers can use the same model to address multiple issues, reducing the need for various models. In code generation [83] [79], test case generation [140] [142], and software security [167] [159], there are instances of using CodeLlama. This model is fine-tuned and optimized based on the LLaMA architecture. At its release, it was considered one of the state-of-the-art models for code generation and understanding tasks, showing strong performance and potential compared to other models like Codex. Another potential reason is the previous successful applications and research outcomes that have proven these models’ effectiveness, further enhancing researchers’ trust and reliance on them. Compared to models that perform well in specific domains, in intelligent agent development, there is a preference for using general-purpose large models to ensure that the core of the agent possesses excellent text comprehension abilities, allowing for further reasoning, planning, and task execution. From the Figure.10, we can also observe that research in the code generation and software development fields adopts a wide variety of models, further indicating the extensive attention this area receives and the excellent performance of models in code generation task.
Figure 11: Distribution of LLMs and Agents across six topics
CODE REQ AUTO DESIGN SEC TEST
CODE X 1 0 2 3 1
REQ 1 X 1 0 2 0
AUTO 0 1 X 6 5 1
DESIGN 2 0 6 X 1 0
SEC 3 2 5 1 X 2
TEST 1 0 1 0 2 X
TABLE XI: Overlap of Papers Among Different Topics (REQ DENOTES ”Requirement Engineering and Documentation”, CODE DENOTES ”Code Generation and Software Development”, AUTO DENOTES ”Autonomous Learning and Decision Making”, DES DENOTES ”Software Design and Decision Making”, SEC DENOTES ”Software Security and Maintenance”)
@@ X-B Topics Overlapping
Figure.11 shows the distribution of all collected literature across six themes. For LLM-type literature, the theme of software security and maintenance accounts for nearly 30%, whereas test case generation accounts for less than 10%. This trend is similarly reflected in the LLM-based agent literature. Research on using LLM-based agents to address requirements engineering and test case generation is relatively sparse. Requirements engineering is a new endeavor for LLM-based agents, and using the entire agent framework to generate test cases might be considered excessive. Therefore, more research tends to evaluate and explore the changes LLMs bring within the agent framework, such as autonomous decision-making abilities and capabilities in software maintenance and repair.
Table‘XI presents the number of papers spanning multiple themes. For instance, five papers can be classified under both software security and maintenance and autonomous learning and decision making. These two themes also overlap the most with other themes, indicating that LLMs and LLM-based agent research is broad and these tasks often require integrating knowledge and techniques from various fields such as code generation, design, and testing. The significant overlap reflects the close interrelation between these themes and other areas. For example, autonomous learning and decision-making often involve the model’s ability to autonomously learn and optimize decision trees, techniques that are applied in many specific software engineering tasks. Similarly, software security and maintenance typically require a combination of multiple techniques to enhance security, such as automatic code generation tools and automated testing frameworks [71] [80] [83] [102]. The overlap in literature highlights the increasing need for integrating methods and techniques from different research areas within software engineering. For instance, ensuring software security relies not only on security measures but also on leveraging code generation, automated testing, and design optimization technologies. Similarly, autonomous learning and decision-making require a comprehensive consideration of requirements engineering, code generation, and system design. Moreover, it suggests that certain technologies and methods possess strong commonality. For instance, LLM-based agents enhance capabilities in code generation, test automation, and security analysis through autonomous learning and decision-making. This sharing of technologies promotes knowledge exchange and technological dissemination across various fields within software engineering.
@@ X-C Benchmarks and Metrics
As shown in Figure.12, it include the distribution of common benchmarks across six topics. In reality, the number of benchmark datasets used is far greater than what is shown in the figure. Different software engineering tasks use various benchmark datasets for evaluation and testing. For instance, in requirements engineering, researchers often collect user stories or requirement specifications as datasets [55] [63], and these datasets are not well-known public datasets, so they were not included in the statistics. Alternatively, some studies specify their datasets as “Customized GitHub datasets” [168]. Therefore, the benchmark datasets shown in the figure represent commonly used public datasets. For example, MBPP and HumanEval, which have been introduced in previous sections, are frequently used. We can also observe that the datasets used in LLM and LLM-based agents tasks, apart from common public datasets, are different.
For instance, the FEVER [^7] https://fever.ai/dataset/fever.html dataset is often used in agent-related research. In [35], the FEVER dataset is used to test the ExpeL agent’s performance in fact verification tasks. Similarly, the HotpotQA [^8] https://hotpotqa.github.io/ dataset is frequently used in agent-related research for knowledge-intensive reasoning and question-answering tasks. When handling vulnerability repair tasks, LLMs often use the Defects4J [^9] https://github.com/rjust/defects4j benchmark dataset. This dataset contains 835 real-world defects from multiple open-source Java projects, categorized into buggy versions and repaired versions, typically used to evaluate the effectiveness of automated program repair techniques. Despite its extensive use in LLM research, Defects4J is relatively less used in LLM-based agents research. We speculate that this may be because Defects4J primarily evaluates single code repair tasks, which do not fully align with the multi-task and real-time requirements of LLM-based agents. Additionally, new datasets like ConDefects have been introduced [142], focusing on addressing data leakage issues and providing more comprehensive defect localization and repair evaluations.
As shown in Figure.13, it includes the top ten evaluation metrics for LLMs and LLM-based agents. The analysis reveals that the evaluation methods used by both are almost identical. In previous sections, we also discussed that for agents, it is necessary to consider time and computational resource consumption, which is evident from the pie chart. Meanwhile, many studies focus on the code generation capabilities of LLMs, so more evaluation metrics pertain to the correctness and Exact Match of the generated code [73] [69] [30], but overall, the evaluation metrics for LLMs and LLM-based agents in software engineering applications are quite similar.

### v2

@@ IX-C Analysis
Figure 9: Illustration of Comparison Framework Between LLM-based Agent [51] and LLM [183] in Software Security and Maintenance.
@@ IX-E Evaluation Metrics
The evaluation metrics for the llm in the software security and maintenance are quite diverse. Researchers need to consider various factors such as coverage, efficiency, and reliability of the model or framework. Evaluation Metrics like success rate and pass rate are directly related to the performance of LLMs in different scenarios. In Table XI, common standards such as success rate and change rate are frequently used to evaluate the robustness of models when faced with diverse inputs. Time overhead and query number are used to assess the efficiency and resource consumption of models when performing specific tasks. Additionally, ROC AUC, F1 score, and accuracy are important for evaluating the model’s ability to identify vulnerabilities, especially in binary classification tasks. In code repair tasks, metrics such as compilability and plausibility are very common, these metrics ensure that the generated solutions are correct and deployable. Common standards like BLEU and CodeBLEU are used to evaluate the quality and human-likeness of generated code, which helps determine if the model’s capabilities and performance are comparable to human performance. Furthermore, domain-specific metrics like tree edit distance and test pass rate are used to evaluate the effectiveness of LLM applications in specialized fields of software engineering, these metrics are used to address the limitations posed by software security and maintenance. In contrast, while LLM-based agents use evaluation metrics similar to those used for LLMs, such as success rate, they also incorporate more subjective metrics for evaluation. These include appropriateness, relevance, and adequacy, which are human-judged standards. Overall, the evaluation metrics used by agents tend to be simpler and easier to understand than those used for LLMs. This is likely because agents handle high-level tasks, such as the success rate of generating potential vulnerabilities and the frequency of agents calling external tools, so they also need to consider computational and time overheads of the overall architecture.
@@
TABLE XI: Evaluation Metrics in Software Security and Maintenance
@@
EvalGPTFix
Number of Correctly Fixed Bugs,
Fix Rate per Prompt Type,
Success Rate over All Test Cases
No
[181]
ChatPHP-DB (2,500 PHP websites)
Vulnerable Parameters (SQLi, XSS, File Upload).
Dynamic + Static + Manual Analysis.
Exploitability Rate, Prepared Statement Ratio.
CWE-based Classification.
No
[186]
@@
MITRE ATTCK framework
@@ 13 Smart Contract Bugs
263 real smart contract vulnerabilities
F1 score, Accuracy, Consistency,Precision, Recall
@@
15 Vulnerabilities from CVE Lib Success Rate, Cost Yes
[193]
SWE-bench
Resolved Ratio, Planning Accuracy,
Line-level Localization Accuracy
Yes
[192]
SWE-bench Lite
Resolved Rate, Pass@1, Pass@3,
Correctness Rate, Average Time per Task,
Token Cost
Yes
@@ X-A Experiment Models
In Sections III-IX, we reviewed and introduced the research on LLMs and LLM-based agent applications in software engineering in recent years. These studies have different research directions and we divided them into six subtopics for classification and discussion. With the advancement of large language models, to better understand their application in various fields, we summarized a total of 110 papers, focusing primarily on discussing the frequency of LLM usage in the field of software engineering
@@
Based on the review of 139 papers, our primary focus is on the models or frameworks utilized by the authors in their experiments. This is due to the fact that these papers often include tests of model performance in specific domains, such as evaluating LLaMA’s performance in code generation. Therefore, during our data collection process, we also included models used for comparison purposes, as these models often represent the state-of-the-art capabilities in their respective fields at the time of the study. In summary, across the 139 papers, we identified a total of 82 unique large language models. We visualized the frequency of these model names in a word cloud for a more intuitive representation, as shown in Figure.10. From the figure, we can observe that models such as GPT-3.5, GPT-4, LLaMA2, and Codex are frequently used. Although close source LLMs cannot be locally deployed or further trained, their exceptional capabilities make them a popular choice for comparison in experiments or for data augmentation, where GPT-4 is used to generate additional data to support the research model frameworks. For instance, researchers might use OpenAI’s API to generate initial text and then employ locally deployed models for further processing and optimization [101, 153, 148, 143].
@@
Therefore, it is not difficult to see that the use of general large models with superior performance to assist development or as a measurement standard has been increasingly used in the vertical field of software engineering in the past two years. In addition, for some fields that have never been touched by LLMs before, many researchers first refer to the model ChatGPT and conduct various performance experiments on the newer GPT-4 [70, 73, 75]. Those models can be integrated into larger systems and combining with other machine learning models and tools, these models can be used to generate natural language responses, while another model handles intent recognition and dialogue management.
Figure 10: Experiment Models Usage Word Cloud.
@@
Although the word cloud provides a rough overview of model usage frequency, it lacks detailed information. To gain deeper insights, we combined a grouped bar chart and a stacked bar chart to further analyze the usage of models in studies across different subtopics. The corresponding bar charts are presented in Figure.11. During the analysis, we found that a large number of models appeared only once. Including these in the bar chart would have made the overall representation cluttered. Therefore, we excluded models that appeared only once and focused on the versatility of the remaining models. On the left side of each subtopic, we depict the models used in LLM-related studies, with the models used in LLM-based agent-related studies highlighted in red-bordered bars. From the figure, it is evident that in the Autonomous Learning and Decision Making subtopic, the number of models used in LLM-based agent-related studies is quite high. Specifically, GPT-4 and GPT-3.5 were used in 12 out of 23 papers and 16 out of 23 papers, respectively. In this subtopic, studies commonly utilized GPT-3.5/4 and LLaMA-2 for research and evaluation. During our analysis, we found that many studies on LLM-based agents evaluated the agents’ ability to mimic human behavior and decision-making or perform some reasoning tasks [131, 141, 137]. Since these studies do not require local deployment, they mainly assess the performance of state-of-the-art models in specific directions, leading to the frequent use of the GPT-family models. Frameworks like [59, 41] constructed LLM-based agents by calling the GPT-4 API, using verbal reinforcement to help language agents learn from their mistakes. Due to the limitations of GPT models, many studies also used LLaMA as the LLM for agents, fine-tuning it on the generated datasets to evaluate the emergence of knowledge and capabilities. Overall, we found that in the Autonomous Learning and Decision Making subtopic, LLM-based agents often use multiple models for testing and performance evaluation in a single task, this results in a significantly higher model usage frequency in this topic compared to others.
@@
Figure 11: Experiment Models Usage in Different Subtopics. (REQ DENOTES ”Requirement Engineering and Documentation”, CODE DENOTES ”Code Generation and Software Development”, AUTO DENOTES ”Autonomous Learning and Decision Making”, DES DENOTES ”Software Design and Decision Making”, SEC DENOTES ”Software Security and Maintenance”)
@@
Therefore, researchers and developers can use the same model to address multiple issues, reducing the need for various models. In code generation [111, 57], test case generation [168, 170], and software security [198, 190], there are instances of using CodeLlama. This model is fine-tuned and optimized based on the LLaMA architecture. At its release, it was considered one of the state-of-the-art models for code generation and understanding tasks, showing strong performance and potential compared to other models like Codex. Another potential reason is the previous successful applications and research outcomes that have proven these models’ effectiveness, further enhancing researchers’ trust and reliance on them. Compared to models that perform well in specific domains, in intelligent agent development, there is a preference for using general-purpose large models to ensure that the core of the agent possesses excellent text comprehension abilities, allowing for further reasoning, planning, and task execution. From the Figure.11, we can also observe that research in the code generation and software development fields adopts a wide variety of models, further indicating the extensive attention this area receives and the excellent performance of models in code generation task.
Figure 12: Distribution of LLMs and Agents across six topics.
Topic A Topic B Overlap Count
REQ CODE 1
REQ SEC 1
CODE TEST 3
CODE AUTO 3
AUTO SEC 3
TEST SEC 1
TABLE XII: Selected Topic Overlaps: Number of Papers Addressing Two Topics Simultaneously (REQ = Requirement Engineering and Documentation, CODE = Code Generation and Software Development, AUTO = Autonomous Learning and Decision Making, SEC = Software Security and Maintenance, TEST = Software Testing and Evaluation)
@@ X-B Topics Overlapping
Figure 12 shows the distribution of all collected literature across six themes. For LLM-type literature, the most represented theme is requirement engineering and documentation, followed closely by software security and maintenance, which together account for nearly half of all LLM papers. In contrast, themes such as test case generation and autonomous decision making are less emphasized. This trend shifts significantly in the LLM-based agent literature. The most prominent theme is autonomous learning and decision making, comprising nearly 30% of all agent-related papers, followed by code generation and development. In comparison, research on using LLM-based agents for requirement engineering and test case generation remains relatively limited, this reflects a tendency to leverage agent frameworks primarily for tasks involving autonomy, adaptability, and iterative decision making, rather than for static tasks such as test generation. As such, more research efforts focus on understanding how agentified LLMs can enhance capabilities in dynamic domains like autonomous decision making and software maintenance.
Table XII presents the selected overlap of papers spanning multiple key themes. Notably, the most frequent overlaps are observed between code generation and software testing [97, 108, 107], as well as between code generation and autonomous learning [110, 114, 54]. These overlaps suggest that tasks in code generation often co-occur with evaluation and learning capabilities, highlighting the integrative nature of LLM and LLM-based agent applications. Additionally, autonomous learning also shares a substantial intersection with software security and maintenance [136, 137, 138], reflecting the importance of adaptive and learning-driven approaches in security-related contexts.
Other overlaps, such as those between requirements engineering and code generation (1 paper), and between testing and security (1 paper), show that even traditionally distinct areas are increasingly interwoven in LLM-related research. These interconnections indicate that achieving high performance in complex software engineering tasks often involves combining methodologies from multiple domains. For example, autonomous learning techniques may be utilized not only in optimizing decision-making processes but also in dynamically generating code or performing real-time security assessments. Similarly, enhancing software security often relies on leveraging capabilities from code generation and testing frameworks [92, 27, 111, 130]. This interconnectedness reinforces the trend that LLM-based agents act as bridges across domains, promoting knowledge transfer and technique reuse within the broader landscape of software engineering.
@@ X-C Benchmarks and Metrics
X-C1 Benchmarks
As shown in Figure 13, we illustrate the distribution of representative benchmark datasets across six major software engineering topics. It is important to note that the number of benchmark datasets used in practice is substantially greater than what is visualized in the figure. This is because many research studies rely on custom or domain-specific datasets that are not widely disseminated or publicly available. For instance, in requirements engineering, researchers often construct datasets from real-world requirement specifications or user stories [70, 81], which are typically not standardized nor openly accessible. Similarly, some studies refer vaguely to their datasets as “Customized GitHub Datasets” [199], thereby excluding them from cross-paper quantitative statistics.
Therefore, the benchmark datasets shown in the figure primarily represent commonly-used public datasets across LLM and LLM-based agent tasks. Among these, HumanEval and MBPP are the most widely adopted across both LLMs and agents for code generation evaluation, given their structured format and popularity in prior benchmark studies. Interestingly, although some overlap exists, the benchmark usage patterns of LLMs and LLM-based agents diverge significantly in focus and diversit. For example, agent-related studies frequently adopt benchmarks that support complex, interactive, or knowledge-intensive reasoning tasks. The FEVER [^9] https://fever.ai/dataset/fever.html and HotpotQA [^10] https://hotpotqa.github.io/ datasets are widely used in agent-based research for fact verification and multi-hop reasoning tasks, respectively. As seen in [40], the FEVER dataset is employed to evaluate the factual consistency of responses generated by the ExpeL agent, while HotpotQA serves as a foundation for question-answering agents requiring long-range dependency tracking and contextual inference.
In the context of software security and maintenance, Defects4J [^11] https://github.com/rjust/defects4j remains a staple benchmark for evaluating code repair capabilities of LLMs. It consists of 835 real-world bugs drawn from multiple open-source Java projects and remains a standard for evaluating the precision and effectiveness of automated program repair. However, it is observed to be rarely used by LLM-based agent studies, potentially due to its focus on single-step repair evaluation, which is less aligned with the multi-turn, tool-augmented, and real-time characteristics of agentic systems. Instead, agent-based studies increasingly turn to more comprehensive benchmarks such as SWE-bench, ProjectDev, or API-Bank, which emphasize multi-stage workflows and tool invocation.
Emerging benchmarks like ConDefects [170] and EvalPlus are designed to mitigate data leakage and improve real-world representativeness. They also provide fine-grained evaluation for sub-tasks like defect localization, which could be more suitable for future multi-agent testing scenarios.
X-C2 Evaluation Metrics.
As shown in Figure 14, we summarize the top 10 evaluation metrics most frequently used in LLM and LLM-based agent studies. The analysis reveals both convergence and divergence in metric preferences between the two paradigms. Traditional LLM research continues to emphasize structural correctness and task-specific quality, with top metrics including Accuracy (20.9%), F1 Score (11.9%), Pass Rate (16.4%), Correctness (11.9%), and Exact Match (6.0%). These metrics are particularly common in structured tasks such as classification [94], code generation [90], and bug repair [35].
In contrast, LLM-based agent research introduces broader, more dynamic evaluation dimensions. The most frequent metric used by agents is Success Rate (20.0%), followed by Pass@1 (12.9%), Accuracy (15.7%), and Recall/Precision/F1 Score (each 8.6%). Furthermore, agent-based studies introduce agent-specific metrics such as Efficiency (5.7%), Cost (4.3%), Correctness Rate (4.3%), and Win Rate (4.5%), which reflect not only task completion, but also resource consumption and comparative agent performance.
It is noteworthy that although both paradigms rely on overlapping core metrics, their usage context differs. For instance, while Accuracy remains the dominant metric for both LLMs and agents too access many software task, it is often interpreted differently static classification accuracy for LLMs versus dynamic subtask completion accuracy for agents [139]. The use of Success Rate as the most frequent agent metric illustrates the community’s emphasis on end-to-end, multi-step success in goal-oriented workflows, such as tool usage or environment interaction [40, 55].
Moreover, qualitative and user-centric evaluation metrics like Alignment, Understandability, and Specifiability have emerged in both LLM and agent studies. These metrics are crucial for measuring the model’s interpretability, instructional clarity, and alignment with user goals, especially in interactive multi-turn or development-assistive settings. In summary, while both LLMs and agents share a substantial subset of evaluation metrics, their distinct execution modes, static inference vs. autonomous orchestration, result in different emphases in metric design and interpretation. This trend underscores the need for context-aware evaluation frameworks to assess the next generation of LLM-based agents.
Figure 13: Distribution of Benchmarks.
Figure 14: Top 10 Evaluation Metrics.
X-D Challenges and Opportunities
As demonstrated across the six key software engineering domains covered in this survey, the transition from LLMs to LLM-based Agents not only introduces new capabilities but also brings about novel technical and theoretical challenges. This part consolidates insights from the reviewed literature, summarizing the core challenges faced by LLM-based Agents in software engineering, and explores the new opportunities enabled by this paradigm shift.
1) Lack of standardized agent definitions and evaluation protocols. Although a working definition is provided in Section III, there is currently no consensus in the academic community regarding the level of autonomy, planning capability, and tool usage required for an LLM to qualify as an agent. Existing studies (e.g., [55, 40, 56]) have proposed varying degrees of agent behavior, such as the self-improvement in Reflexion [59] and the reasoning mechanism in Graph-RAG [83], yet a unified evaluation benchmark remains absent. This fragmentation hinders reproducibility, horizontal comparisons, and the development of a cohesive theoretical foundation.
2) Workflow complexity and error propagation in multi-agent systems. While multi-agent collaboration enhances modularity and scalability [58, 43], it also introduces challenges in synchronization, error recovery, and context sharing. As shown in [84, 81], role-based collaboration can improve system performance, yet ensuring state consistency and semantic alignment among agents remains an open problem. Unlike traditional single LLM pipelines, asynchronous interactions between agents are more prone to cascading misunderstandings, especially in long-tail tasks such as requirement elicitation or security analysis.
3) Tool integration bottlenecks and external dependency management. Many LLM-based Agents rely on external tools to complete tasks (e.g., API-Bank in ToolCoder [139], browser environments in AutoCodeRover [55]), but integrating tools into different software engineering workflows (compilers, test frameworks, version control systems) often leads to fragile interfaces and poor adaptability. Moreover, there is a lack of standardized evaluation strategies for measuring agent performance under tool latency, failure, or partial observability.
4) Lack of cross-task generalization and knowledge transfer capabilities. Although some agents exhibit expertise in specific tasks, such as program repair or requirement compliance checking, agents capable of transferring knowledge across tasks remain rare. Frameworks like MAGIS [85] and SWE-agent [53] attempt to accumulate experience through multiple interactions, but broader meta-learning or task-agnostic transfer mechanisms have yet to be established.
5) Data scarcity and insufficient simulation environments constrain agent behavior evaluation. Traditional LLM benchmarks (e.g.HumanEval, MBPP) focus on static input-output tasks, whereas agents in domains such as requirement engineering and software ssecurity, shown in works like [84, 83, 44] require multi-turn, multi-role simulations. The scarcity of such interactive datasets limits our ability to evaluate advanced agent behaviors such as negotiation, role conflict resolution, and tool-driven reasoning.
6) Cognitive transparency and trustworthiness issues. As agent autonomy increases, their reasoning and planning processes become increasingly opaque. Although studies like [69, 65] propose methods aligned with user expectations, the interpretability of decision traces, particularly in critical scenarios such as CoT + ToT planning for security sensitive tasks, it remains a key obstacle to deployment.
X-E Opportunities for Future Research
1) Towards a unified agent taxonomy and evaluation framework to enable end-to-end agent-driven software development pipelines. The research community may establish shared definitions and multidimensional evaluation scales encompassing autonomy, tool usage, context management, and collaborative reasoning. Benchmarks such as CAASD [79] and API-Bank [114] provide a solid foundation for agent-protocol-driven task construction. From requirement elicitation [65, 81], system design [69], testing [107], to secure deployment [172], agent modules can gradually be connected into an autonomous software engineering pipeline. This includes linking pre-development planning agents with post-deployment monitoring agents, forming a learning-enabled, decision-autonomous ecosystem.
2) Enhancing agent memory and reflection mechanisms. As illustrated by LLM-based Agents in this survey [103, 59, 40, 82, 99], agents with explicit memory pools, experience tracking, and self-evaluation outperform stateless systems in complex debugging and requirement construction tasks. Future research may incorporate structured memory, error trajectory analysis, and goal-driven retrospection to enable stronger long-term planning capabilities.
3) Technical collaboration between agents and humans, and embodied agents with multimodal capabilities in software engineering. Several studies [152, 153] emphasize the importance of designing workflows and evaluation processes with human-in-the-loop. Future agents should be able to support role-specific users (e.g., test engineers, product managers), achieving alignment in output granularity, decision traceability, and task handover mechanisms. Inspired by VOYAGER [109, 43, 53], SE agents could be extended to handle graphical user interface operations, visual debugging, and even embedded system control. Such multimodal capabilities will allow agents to break free from purely text-based environments and operate in richer SE practice settings.
4) Practical deployment of LLM-based Agents in software development. Most current LLM-based Agents remain focused on function-level or module-level code completion, falling short in real-world project development tasks that require complex dependency management, version control, and multi-file structure organization [200]. For LLM-based Agents to transition into practical engineering, they must possess project-level context modeling and dynamic tracking capabilities. For example, the Dynamic Code Graph Generator (DCGG) module [108] continuously models cross-file dependencies, class structures, and component versions to enhance adaptability to large-scale system construction tasks. Moreover, agents should evolve from being single-execution actuators to “development workflow collaborators” with role awareness. Frameworks like MapCoder [103], MetaGPT [117], and ClarifyGPT [115] demonstrate diversified, role-specific agent task division mechanisms that significantly improve performance on complex tasks. Furthermore, frameworks such as AgentCoder and L2MAC emphasize the importance of task-decomposed and role-distributed agent collaboration as a key enabler for transitioning LLM-based Agents from static completion to real-world software engineering practice. Compared to traditional single-model or static scripting pipelines, such multi-agent architectures offer greater robustness and task adaptability, especially demonstrating superior scalability and debugging traceability in project-level engineering.